# Stability of quasi-particle creation and multiband geometry in fractional Chern insulators under magnetic fields

### v1

@@ IV.2 Kapit-Mueller model
Figure 4: One-plaquette Chern number $\mathcal{C}$ in the Kapit-Mueller model with $U=1.0,N=4$ at (a) $N_{\phi}=+1,V_{p}=-1.0$ and (b) $N_{\phi}=-1,V_{p}=+1.0$. (c) Variance $\mathcal{V}[\mathcal{C}]$ for $U=1\sim 8$ and the flux density $N_{\phi}=0,\pm 1$. The yellow region with large $\mathcal{V}[\mathcal{C}]$ corresponds to the non-FCI state around the non-interacting limit $U=0$.
@@
We first discuss the quasi-hole creation in the Kapit-Mueller model. In Fig. 3, we show spectral-flow for the Kapit-Mueller model with one additional positive flux quantum $N_{\phi}=1$. The counting-rule for the quasi-holes is #$\#_{\text{qh}}=13$ for the present system with $N=4,N_{s}/3=12,C_{\text{s}}N_{\phi}=1$, and correspondingly the lowest 13 energy levels are shown in blue in Fig. 3(a). There exists a clear energy gap between the set of the blue curves and others, which means that the counting-rule is satisfied. Figure 3(b) represents spectral flow for the system with the pinning potential $V_{p}=1.0$, where the blue curves represent the three lowest energy states. Now these three fold quasi-degenerate states are well separated from other states, and they are expected to exhibit the fractionally quantized Hall conductance as mentioned in the previous section [35]. Indeed, we find that the many-body Chern number for the three fold quasi-degenerate ground states for $V_{p}=1.0$ is $C=1$ corresponding to the $\nu=1/3$ FQHE. In addition, the one-plaquette Chern number $\mathcal{C}(\vec{\theta})$ is nearly flat in the parameter space as shown in Fig. 4(a). These results are fully consistent with the qualitative discussions in the previous section. Namely, the quasi-holes obeying the counting rule are created by the magnetic field and then get localized by the pinning potential, which consequently leads to the FCI ground state for $\phi\neq 0$ and $V_{p}\neq 0$. Qualitatively same results are obtained for other values of the interaction strength, $U=1.0\sim 8.0$. The FCI state is unstable and a non-FCI state is realized only around the non-interacting limit $U=0$.
@@
Next we consider the quasi-electron creation in the presence of one additional negative flux quantum $N_{\phi}=-1$. Spectral flow with no pinning potential at $N=4,U=1.0$ is shown in Fig. 3(c). As mentioned in the previous section, the counting rule for the quasi-electrons is #$\#_{\text{qe}}=11$. It is found that there are 11 fold quasi-degeneracy for the present Kapit-Mueller model, and thus the quasi-electrons are stably created by the mangnetic field. Similarly to the quasi-hole case, spectral flow with the pinning potential (Fig. 3(d)) also shows the three fold quasi-degenerate ground states which are separated from other excited states with a gap. Note that, in the quasi-electron case, a negative pinning potential $V_{p}=-1.0$ is necessary to localize the quasi-electron, since it has a positive fractional charge as opposed to the quasi-holes. The many-body Chern number of the three fold quasi-degenerate ground states is found to be $C=1$, and the one-plaquette Chern number is uniform as shown in Fig. 4 (b). Similar results are obtained for $U=1.0\sim 8.0$.
@@ IV.3 Checkerboard model
We present numerical results of the spectral flow for the quasi-hole creation in the checkerboard model at $N=6,U=1.0$ in Fig. 5(a). The blue curves in Fig.5(a) represent the energy levels expected from the counting rule, #$\#_{\text{qh}}=19$ for $N=6,N_{s}/2=18,C_{s}N_{\phi}=1$. However, at $U=1.0$, there is no clear energy gap between the blue curves and others, and it is difficult to identify the ground state sector in contrast to the Kapit-Mueller model in the previous section. Consequently, there are no clear three fold quasi-degenerate ground states with the pinning potential as shown in Fig. 5(b). Note that the pinning potential is sufficiently large $V_{p}=10$ so that density distribution is well localized around the site $j_{0}$, and we have confirmed the absence of the quasi-degeneracy for other values of the pinning potential. This implies that quasi-holes are not created and the underlying FCI state is no longer stable in the presence of the magnetic field. Indeed, the one-plaquette Chern number for the three lowest energy states in Fig. 5(b) diverges and has large fluctuations (Fig. 6(a)). Since the one-plaquette Chern number is nearly flat in the $\theta$-space in a general gapped system in the thermodynamic limit [55, 56], the strong divergence implies that the present system with $U=1$ does not have the three fold topological quasi-degeneracy. Therefore, the ground state is a non-FCI state.
@@
Figure 6: One-plaquette Chern number $\mathcal{C}$ for the three lowest energy states of $\tilde{H}$ in the checkerboard model at (a) $N=6,N_{\phi}=-1,U=1$, (b) $N=6,N_{\phi}=-1,U=2$, (c) $N=6,N_{\phi}=-1,U=3$, (d) $N=6,N_{\phi}=+1,U=1$.
@@
However, as we increase the interaction strength from $U=1.0$ to $U=2.0,3.0$, an energy gap develops in the spectral flow at $V_{p}=0$ and the low energy 19 states become separated from others as seen in Fig. 5(c),(e). Furthermore, in the presence of the pinning potential, we find that there exists a clear gap between the three lowest energy levels and other excitations. The many-body Chern number for these three fold quasi-degenerate states is $C=-1$ and the variance $\mathcal{V}[\mathcal{C}]$ for $U=2.0\sim 3.0$ is suppressed as shown in Figs. 6 and  7. These results suggest that there is a quantum phase transition from the non-FCI state to the stable FCI state as $U$ increases. The phase transition point is located around $U=2.0$, where the diverging behaviors of the variance $\mathcal{V}[\mathcal{C}]$ get suppressed and it becomes rather flat as seen in Fig. 7(b). To be more precise, a quantum phase transition is well defined only in the thermodynamic limit. In a thermodynamically large system, the number of the fluxes is $|N_{\phi}|\gg 1$ for a fixed magnetic field $\phi$ and there will be $|N_{\phi}|$ quasi-particles localized by a pinning potential corresponding to disorder. We have simply assumed that effects of interactions between the pinned quasi-particles are negligible. This enables us to discuss the phase transition based on the exact diagonalization for the small system with $|N_{\phi}|=1$.
Similar phase transitions are found for the systems with different numbers of the fermions, $N=4,5$. The applied flux per plaquette is $\phi=-1/N_{s}\simeq-0.042$ for $N=4$ and $\phi\simeq-0.033$ for $N=5$, which enables us to discuss magnetic field dependence of the quantum phase transition. Similarly to the $N=6$ case, as $U$ increases, energy gaps develop between the #$\#_{\text{qh}}$ lowest energy states and other states at $V_{p}=0$, and also between the three lowest energies and excited states for $V_{p}\neq 0$ (see Appendix B for details). We find that the strength of the interactions $U_{c}$ required to stabilize the FCI increases as $N$ decreases. Besides, the variance of the Chern number is large for small interactions as seen in Fig. 7(b). Based on these numerical results, we suppose that the phase transition occurs at $U_{c}=2.0\sim 3.0$ for $N=5$ and $U_{c}=4.0\sim 5.0$ for $N=4$. This in turn implies that the critical interaction $U_{c}$ increases as the negative magnetic field increases. There would be some finite size effects for smaller system sizes, but we can confirm that it is not relevant in the present system. First, when the system size is changed, quasi-holes are stably created for $N=5,U=1.0$ at $\phi=0$ (Appendix A), which means that the absence of the FCI state for $N=5,U=1.0$ at $\phi\simeq-0.033$ is due to the magnetic field. Second, as will be discussed in the following, the FCI state becomes stable when the direction of the magnetic field is changed, $\phi\to-\phi$. Therefore, the stability of the FCI state is dominated by the magnetic field.
@@
Figure 7: (a) Many-body Chern number $C$ for various system sizes and fluxes. (b) Variance $\mathcal{V}[\mathcal{C}]$. The inset is an enlarged view so that the diverging behaviors are clearly seen.
@@
In contrast to the quasi-hole case, creation of the quasi-electrons is stable even for small interactions. As shown in Fig. 8 and Fig. 15 in Appendix B, the counting rule derived from the numerical results for the FQHE and the three fold quasi-degeneracy are realized for $N=4,5,6$ at $U\approx 1.0$. Moreover, the many-body Chern number $C=-1$ is robust even at the small interaction $U=1.0$ for all $N=4,5,6$. Note that the variance ${\mathcal{V}}[C]$ is slightly large at $U=1.0$ for $N=4$ and $5$, but this may be due to finite size effects and ${\mathcal{V}}[C]$ is quickly suppressed for larger $U$. Thus, the stable FCI region is extended over a wide range of interaction strength and magnetic fields for the quasi-electron creation.
In Fig. 9, we summarize the numerical results of the variance ${\mathcal{V}}[C]$ as an indicator for the FCI state. The many-body Chern number for the three fold quasi-degenerate ground states with $V_{p}\neq 0$ is $C=-1$ and it has small fluctuation ${\mathcal{V}}[C]$ in the FCI state, while the Chern number is not well defined and the variance diverges in the non-FCI state. The flux $|\phi|=1/N_{s}\simeq 0.042,0.033,0.028$ corresponds to the different numbers of the fermions, $N=4,5,6$, respectively. We see that the creation of the quasi-particles and the underlying FCI state are stable for a wide range of the parameters, while the non-FCI state is realized around the non-interaction region $U=0$ and under negative fluxes even with moderate interactions. Based on these results, we obtain the expected phase diagram Fig. 2(b) in Sec. III. It is difficult to clarify details of the non-FCI state within the present calculations, and further discussions are left for a future study.
Figure 9: Variance $\mathcal{V}[\mathcal{C}]$ in the checkerboard model as a function of the flux $\phi$ and the interaction $U$. The many-body Chern nubmer is $C=-1$ in the region with $\mathcal{V}[\mathcal{C}]\simeq 0$, while it is not well-defined when $\mathcal{V}[\mathcal{C}]$ is not negligibly small. The variance in the white-colored region cannot be interpolated from that in neighboring regions.
@@ V MultiBand Quantum Geometry
where $\mathcal{P}(\bm{k})=\sum_{a=1}^{N_{\text{b}}}\ket{u^{a}_{\bm{k}}}\bra{u^{a}_{\bm{k}}}$ | is the projection operator onto the target $N_{\text{b}}$ bands which we focus on. The multiband quantum metric and Berry curvature are given as the symmetric and anti-symmetric parts of $\eta$,
@@
where tr is a trace over the indices $\alpha,\beta$. Since the multiband quantum geometry plays fundamental roles for realization of the FCI state, we naively expect that the multiband quantum geometry under magnetic fields is relevant also for stability of the quasi-particle creation. More precisely, our expectation is that if a clean system with a “commensurate” filling (i.e. $\nu=1/3$) has good multiband quantum geometry under a magnetic field $\phi\neq 0$ and thus exhibits an FCI ground state, creation of spatially extended quasi-particles induced by the “incommensurate” filling (i.e. $\nu=1/3+\delta\nu$ according to the Streda’s formula) is stable at the same magnetic field $\phi$. The itinerant quasi-particles will be localized if there is a pinning potential. It turns out that this conjecture is indeed helpful for understanding the exact diagonalization results from the perspective of band properties.
@@
We show the width $W$ of the multiband in Fig. 11, which is defined as the difference between the maximum and minimum energy values within the multiband. In the Kapit-Mueller model, the bandwidth $W$ is nearly unchanged for $\phi<0$ and its change is non-zero but small for $\phi>0$. In the checkerboard model, the bandwidth varies largely at $\phi\neq 0$. The negative magnetic field which induces quasi-hole creation enhances the multiband width more strongly than the positive magnetic field corresponding to the quasi-electron creation. This indicates that a larger interaction is required to realize the multiband FCI for the negative strong magnetic field. However, within the range of $\phi$ considered in the exact diagonalization (the dashed vertical lines in Fig. 11), the difference of the bandwidth for ±$\pm\phi$ is relatively small. It would not be relevant for the asymmetric stability of the FCI under $\phi>0$ and $\phi<0$ found in the exact diagonalization.
On the other hand, the quantum geometry is shows characteristic behaviors depending on the field direction in the checkerboard model and hence is useful for understanding the stability of the quasi-particle creation, as discussed below. The breakdown of the trace condition can be characterized by the following quantity introduced in the previous study [16],
$T[\eta]\coloneqq\text{Tr}[\text{tr}g]-\text{Tr}[\mathcal{F}].$ (21)
We show $T[\eta]$ for the two models in Fig. 12 which are calculated for the magnetic Brillouin zones of sufficiently large systems. In the Kapit-Mueller model, the trace condition is almost satisfied for all $|\phi|\leq 0.1$. The band properties favor the multiband FCI under magnetic fields. This is consistent with the stable quasi-particle creation by magnetic fields in the Kapit-Mueller model as seen in the exact diagonalization in the previous section. Besides, the small $T[\eta]$ implies that the FCI phase is extended to a moderately large $|\phi|$.
@@
In the checkerboard model, $T[\eta]$ decreases with the addition of positive magnetic fields $\phi>0$, whereas they increase for $\phi<0$. The multiband trace condition $T[\eta]=0$ is almost fully achieved around $\phi=0.08\sim 0.10$. This implies enhanced stability of the multiband FCI state under $\phi>0$, although the bandwidth $W$ increases to some extent. On the other hand, $T[\eta]$ significantly increases for $\phi<0$. These band properties suggest that the FCI at $\phi<0$ is less stable than that at $\phi>0$, which is consistent with the behaviors of the quasi-particle creation discussed in the previous section. In this sense, the stability of the quasi-particle creation by magnetic fields can be understood based on the multiband quantum geometry. Note that the relation between the quasi-particle creation and the multiband quantum geometry may not be valid for a clean system under strong magnetic fields, where there is a finite density of quasi-particles and interactions between them. Stability of quasi-particles with a finite density would be an interesting future problem.

### v2

@@ IV.2 Kapit-Mueller model
Figure 4: One-plaquette Chern number $\mathcal{C}$ in the Kapit-Mueller model with $U=1.0,N=4$ at (a) $N_{\phi}=+1,V_{p}=-1.0$ and (b) $N_{\phi}=-1,V_{p}=+1.0$. (c) Variance $\mathcal{V}[\mathcal{C}]$ for ${|V_{p}|=1,}U=1\sim 8$ and the flux density $N_{\phi}=0,\pm 1$. The yellow region with large $\mathcal{V}[\mathcal{C}]$ corresponds to the non-FCI state around the non-interacting limit $U=0$.
@@
We first discuss the quasihole creation in the Kapit-Mueller model. In Fig. 3, we show spectral-flow as a function of $\theta_{y}$ for the Kapit-Mueller model with one additional positive flux quantum $N_{\phi}=1$. Essentially the same results can be obtained also as a function of $\theta_{x}$, since the boundary twisting has only negligibly small effects in a gapped system in the thermodynamic limit [62]. The counting-rule for the quasiholes is #$\#_{\text{qh}}=13$ for the present system with $N=4,N_{s}/3=12,C_{\text{s}}N_{\phi}=1$, and correspondingly the lowest 13 energy levels are shown in blue in Fig. 3(a). There exists a clear energy gap between the set of the blue curves and others, which means that the counting-rule is satisfied. Figure 3(b) represents spectral flow for the system with the pinning potential $V_{p}=1.0$, where the blue curves represent the three lowest energy states. Now these three fold quasi-degenerate states are well separated from other states, and they are expected to exhibit the fractionally quantized Hall conductance as mentioned in the previous section [39, 40]. Indeed, we find that the many-body Chern number for the three fold quasi-degenerate ground states for $V_{p}=1.0$ is $C=1$ corresponding to the $\nu=1/3$ FQHE. In addition, the one-plaquette Chern number $\mathcal{C}(\vec{\theta})$ is nearly flat in the parameter space as shown in Fig. 4(a). These results are fully consistent with the qualitative discussions in the previous section. Namely, the quasiholes obeying the counting rule are created by the magnetic field and then get localized by the pinning potential, which consequently leads to the FCI ground state for $\phi\neq 0$ and $V_{p}\neq 0$. Qualitatively same results are obtained for other values of the interaction strength, $U=1.0\sim 8.0$. The FCI state is unstable and a non-FCI state is realized only around the non-interacting limit $U=0$. We have also performed the same calculations for different values of the pinning potential, $V_{p}=0.1$ and $V_{p}=10$. For the small pinning potential, the energy scale of the quasihole states is smaller than the energy gap between the ground states and the excited states at $V_{p}=0$, for which our projection scheme is well justified. We find that the qualitative behaviors do not change for both $V_{p}=0.1,10$ (not shown). Furthermore, we have checked that essentially the same results can be obtained even when we take the full Hilbert space into account without the projection onto the low energy sector.
@@
Next we consider the quasielectron creation in the presence of one additional negative flux quantum $N_{\phi}=-1$. Spectral flow with no pinning potential at $N=4,U=1.0$ is shown in Fig. 3(c). As mentioned in the previous section, the counting rule for the quasielectrons is #$\#_{\text{qe}}=11$. It is found that there are 11 fold quasi-degeneracy for the present Kapit-Mueller model, and thus the quasielectrons are stably created by the mangnetic field. Similarly to the quasihole case, spectral flow with the pinning potential (Fig. 3(d)) also shows the three fold quasi-degenerate ground states which are separated from other excited states with a gap. Note that, in the quasielectron case, a negative pinning potential $V_{p}=-1.0$ is necessary to localize the quasielectron, since it has a positive fractional charge as opposed to the quasiholes. The many-body Chern number of the three fold quasi-degenerate ground states is found to be $C=1$, and the one-plaquette Chern number is uniform as shown in Fig. 4 (b). Similar results are obtained for $U=1.0\sim 8.0$ and $V_{p}=-0.1,-10$. Calculations without the projection also give essentially same results.
@@ IV.3 Checkerboard model
We present numerical results of the spectral flow for the quasihole creation in the checkerboard model at $N=6,U=1.0$ in Fig. 5(a). The blue curves in Fig.5(a) represent the energy levels expected from the counting rule, #$\#_{\text{qh}}=19$ for $N=6,N_{s}/2=18,C_{s}N_{\phi}=1$. However, at $U=1.0$, there is no clear energy gap between the blue curves and others, and it is difficult to identify the ground state sector in contrast to the Kapit-Mueller model in the previous section. Consequently, there are no clear three fold quasi-degenerate ground states with the pinning potential as shown in Fig. 5(b). Note that the pinning potential is sufficiently large $V_{p}=10$ so that density distribution is well localized around the site $j_{0}$, and we have confirmed the absence of the quasi-degeneracy for other values of the pinning potential, $V_{p}=0.1,1$. This implies that quasiholes are not created and the underlying FCI state is no longer stable in the presence of the magnetic field. Indeed, the one-plaquette Chern number for the three lowest energy states in Fig. 5(b) diverges and has large fluctuations (Fig. 6(a)). Since the one-plaquette Chern number is nearly flat in the $\theta$-space in a general gapped system in the thermodynamic limit [61, 62], the strong divergence implies that the present system with $U=1$ does not have the three fold topological quasi-degeneracy well separated from excited states. Therefore, the ground state is a non-FCI state.
@@
Figure 6: One-plaquette Chern number $\mathcal{C}$ for the three lowest energy states of $\tilde{H}$ in the checkerboard model at (a) $N=6,N_{\phi}=-1,U=1$, (b) $N=6,N_{\phi}=-1,U=2$, (c) $N=6,N_{\phi}=-1,U=3$, (d) $N=6,N_{\phi}=+1,U=1$. The pinning potential is $V_{p}=10$ for the quasiholes and $V_{p}=-1$ for the quasielectrons.
@@
However, as we increase the interaction strength from $U=1.0$ to $U=2.0,3.0$, an energy gap develops in the spectral flow at $V_{p}=0$ and the low energy 19 states become separated from others as seen in Fig. 5(c),(e). This implies that the spatially extended quasihole states may be created by the magnetic field when the interaction is strong. Furthermore, in the presence of the pinning potential, we find that there exists a clear gap between the three lowest energy levels and other excitations. The many-body Chern number for these three fold quasi-degenerate states is $C=-1$ and the variance $\mathcal{V}[\mathcal{C}]$ for $U=2.0\sim 3.0$ is suppressed as shown in Figs. 6 and  7. These results suggest that there is a quantum phase transition from the non-FCI state to the stable FCI state as $U$ increases. The phase transition point is located around $U=2.0$, where the diverging behaviors of the variance $\mathcal{V}[\mathcal{C}]$ get suppressed and it becomes rather flat as seen in Fig. 7(b). To be more precise, a quantum phase transition is well defined only in the thermodynamic limit. In a thermodynamically large system, the number of the fluxes is $|N_{\phi}|\gg 1$ for a fixed magnetic field $\phi$ and there will be $|N_{\phi}|$ quasiparticles localized by a pinning potential corresponding to disorder. We have simply assumed that effects of interactions between the pinned quasiparticles are negligible. This enables us to discuss the phase transition based on the exact diagonalization for the small system with $|N_{\phi}|=1$. The above discussions are based on the results with the strong pinning potential $V_{p}=10$, but we have confirmed that the qualitative results that the FCI state is unstable at weak interactions is unchanged for other values of the pinning potential, $V_{p}=0.1,1$. We have also checked that qualitatively similar results can be obtained without the projection onto the low energy sector. However, it is difficult to quantitatively evaluate the critical interaction strength, and it will be an interesting future problem.
@@
Figure 7: (a) Many-body Chern number $C$ for various system sizes and fluxes. (b) Variance $\mathcal{V}[\mathcal{C}]$. The inset is an enlarged view so that the diverging behaviors are clearly seen. The pinning potential is $V_{p}=10$ for the quasiholes and $V_{p}=-1$ for the quasielectrons.
Similar phase transitions are found for the systems with different numbers of the fermions, $N=4,5$. The applied flux per plaquette is $\phi=-1/N_{s}\simeq-0.042$ for $N=4$ and $\phi\simeq-0.033$ for $N=5$, which enables us to discuss magnetic field dependence of the quantum phase transition. Similarly to the $N=6$ case, as $U$ increases, energy gaps develop between the #$\#_{\text{qh}}$ lowest energy states and other states at $V_{p}=0$, and also between the three lowest energies and excited states for $V_{p}\neq 0$ (see Appendix B for details). We find that the strength of the interactions $U_{c}$ required to stabilize the FCI increases as $N$ decreases. Besides, the variance of the Chern number is large for small interactions as seen in Fig. 7(b). Based on these numerical results, we suppose that the phase transition occurs at $U_{c}=2.0\sim 3.0$ for $N=5$ and $U_{c}=4.0\sim 5.0$ for $N=4$. This in turn implies that the critical interaction $U_{c}$ increases as the negative magnetic field increases. There would be some finite size effects for smaller system sizes, but we can confirm that it is not relevant in the present system. First, when the system size is changed, quasiholes are stably created for $N=5,U=1.0$ at $\phi=0$ (Appendix A), which means that the absence of the FCI state for $N=5,U=1.0$ at $\phi\simeq-0.033$ is due to the magnetic field. Second, as will be discussed in the following, the FCI state becomes stable when the direction of the magnetic field is changed, $\phi\to-\phi$. Therefore, the stability of the FCI state is dominated by the magnetic field.
@@
In contrast to the quasihole case, creation of the quasielectrons is stable even for small interactions. As shown in Fig. 8 and Fig. 15 in Appendix B, the counting rule derived from the numerical results for the FQHE and the three fold quasi-degeneracy are realized for $N=4,5,6$ at $U\approx 1.0$. Moreover, the many-body Chern number $C=-1$ is robust even at the small interaction $U=1.0$ for all $N=4,5,6$. Note that the variance ${\mathcal{V}}[C]$ is slightly large at $U=1.0$ for $N=4$ and $5$, but this may be due to finite size effects and ${\mathcal{V}}[C]$ is quickly suppressed for larger $U$. Thus, the stable FCI region is extended over a wide range of interaction strength and magnetic fields for the quasielectron creation.
Figure 9: Variance $\mathcal{V}[\mathcal{C}]$ in the checkerboard model as a function of the flux $\phi$ and the interaction $U$. The many-body Chern nubmer is $C=-1$ in the region with $\mathcal{V}[\mathcal{C}]\simeq 0$, while it is not well-defined when $\mathcal{V}[\mathcal{C}]$ is not negligibly small. The variance in the white-colored region cannot be interpolated from that in neighboring regions. The pinning potential is $V_{p}=10$ for the quasiholes and $V_{p}=-1$ for the quasielectrons.
In Fig. 9, we summarize the numerical results of the variance ${\mathcal{V}}[C]$ as an indicator for the FCI state. The many-body Chern number for the three fold quasi-degenerate ground states with $V_{p}\neq 0$ is $C=-1$ and it has small fluctuation ${\mathcal{V}}[C]$ in the FCI state, while the Chern number is not well defined and the variance diverges in the non-FCI state. The flux $|\phi|=1/N_{s}\simeq 0.042,0.033,0.028$ corresponds to the different numbers of the fermions, $N=4,5,6$, respectively. We see that the creation of the quasiparticles and the underlying FCI state are stable for a wide range of the parameters, while the non-FCI state is realized around the non-interaction region $U=0$ and under negative fluxes even with moderate interactions. Based on these results, we obtain the expected phase diagram Fig. 2(b) in Sec. III. It is difficult to clarify details of the non-FCI state within the present calculations, and further discussions are left for a future study.
@@ V MultiBand Quantum Geometry
where $\mathcal{P}(\bm{k})=\sum_{a=1}^{N_{\text{b}}}\ket{u^{a}_{\bm{k}}}\bra{u^{a}_{\bm{k}}}$ | is the projection operator onto the target $N_{\text{b}}$ bands which we focus on. The factor $N_{\rm b}A$ is redundant, but it has been introduced for later convenience in Appendix C. The multiband quantum metric and Berry curvature are given as the symmetric and anti-symmetric parts of $\eta$,
@@
where tr is a trace over the indices $\alpha,\beta$ [16, 17, 18]. This is a set of ideal conditions for each constituent single-band. Since the multiband quantum geometry plays fundamental roles for realization of the FCI state, we naively expect that the multiband quantum geometry under magnetic fields is relevant also for stability of the quasiparticle creation. More precisely, our expectation is that if a clean system with a “commensurate” filling (i.e. $\nu=1/3$) has good multiband quantum geometry under a magnetic field $\phi\neq 0$ and thus exhibits an FCI ground state, creation of spatially extended quasiparticles induced by the “incommensurate” filling (i.e. $\nu=1/3+\delta\nu$ according to the Streda’s formula) is stable at the same magnetic field $\phi$. The itinerant quasiparticles will be localized if there is a pinning potential. It turns out that this conjecture is indeed helpful for understanding the exact diagonalization results from the perspective of band properties.
@@
We show the width $W$ of the multiband in Fig. 11, which is defined as the difference between the maximum and minimum energy values within the multiband. In the Kapit-Mueller model, the bandwidth $W$ is nearly unchanged for $\phi<0$ and its change is non-zero but small for $\phi>0$. In the checkerboard model, the bandwidth varies largely at $\phi\neq 0$. The negative magnetic field which induces quasihole creation enhances the multiband width more strongly than the positive magnetic field corresponding to the quasielectron creation. This indicates that a larger interaction is required to realize the multiband FCI for the negative strong magnetic field. However, within the range of $\phi$ considered in the exact diagonalization (the dashed vertical lines in Fig. 11), the difference of the bandwidth for ±$\pm\phi$ is relatively small. It would not be relevant for the asymmetric stability of the FCI under $\phi>0$ and $\phi<0$ found in the exact diagonalization.
On the other hand, the quantum geometry is shows characteristic behaviors depending on the field direction in the checkerboard model and hence is useful for understanding the stability of the quasiparticle creation, as discussed below. The breakdown of the trace condition can be characterized by the following quantity introduced in the previous study [16],
$T[\eta]\coloneqq\text{Tr}[\text{tr}g]-\text{Tr}[\mathcal{F}].$ (21)
We show $T[\eta]$ for the two models in Fig. 12 which are calculated for the magnetic Brillouin zones of sufficiently large systems. In the Kapit-Mueller model, the trace condition is almost satisfied for all $|\phi|\leq 0.1$. The band properties favor the multiband FCI under magnetic fields. This is consistent with the stable quasiparticle creation by magnetic fields in the Kapit-Mueller model as seen in the exact diagonalization in the previous section. Besides, the small $T[\eta]$ implies that the FCI phase is extended to a moderately large $|\phi|$.
@@
In the checkerboard model, $T[\eta]$ decreases with the addition of positive magnetic fields $\phi>0$, whereas they increase for $\phi<0$. The multiband trace condition $T[\eta]=0$ is almost fully achieved around $\phi=0.08\sim 0.10$. This implies enhanced stability of the multiband FCI state under $\phi>0$, although the bandwidth $W$ increases to some extent. On the other hand, $T[\eta]$ significantly increases for $\phi<0$. These band properties suggest that the FCI at $\phi<0$ is less stable than that at $\phi>0$, which is consistent with the behaviors of the quasiparticle creation discussed in the previous section. In this sense, the stability of the quasiparticle creation by magnetic fields can be understood based on the multiband quantum geometry. Note that the relation between the quasiparticle creation and the multiband quantum geometry may not be valid for a clean system under strong magnetic fields, where there is a finite density of quasiparticles and interactions between them. Stability of quasiparticles with a finite density would be an interesting future problem.